OpenStudy (anonymous):

You randomly pick two marbles from a bag containing 3 yellow marbles and 4 yellow marbles and 4 red marbles. you pick the second marble without replacing the first marble. find each probability. 1. p(red then red) 2. p(yellow then red)

OpenStudy (anonymous):

wait so theres 7 yellows?

OpenStudy (anonymous):

no 3 sorry forget the 4 yellow marbles

OpenStudy (anonymous):

haha ok so grabbing a red is 1/7 right? nd no replace ment makes the next marble 1/6 so if i remember correctly you multiply these nnumbers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i could be wrong, im doubting nmyself...

OpenStudy (anonymous):

ima check with a friend...

OpenStudy (anonymous):

stupid friend no answer phone.

OpenStudy (anonymous):

i think your right

OpenStudy (anonymous):

ummm ithink so but doubts

OpenStudy (anonymous):

try

jimthompson5910 (jim_thompson5910):

oops made a typo

OpenStudy (anonymous):

ok @jim_thompson5910 got it right.

jimthompson5910 (jim_thompson5910):

P(red then red) = P(red)*P(red) P(red then red) = (4/7)*(3/6) P(red then red) = 12/42 P(red then red) = 2/7

OpenStudy (anonymous):

the answer is that be cause yu have 4 out of 7 reds, ns once you take out yu now have 3 of the the 6(yu lost one)

jimthompson5910 (jim_thompson5910):

P(yellow then red) = P(yellow)*P(red) P(yellow then red) = (3/7)*(4/6) P(yellow then red) = 12/42 P(yellow then red) = 2/7

OpenStudy (anonymous):

haha wrong answer the first time but i understood it @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so in this case, the probabilities are the same